# Compressing Language Models for Specialized Domains

### v1

@@ 4.2 Domains and Tasks
To assess the efficacy of our approach on downstream tasks, we experiment with what are arguably the two most extensively explored domains in NLP, the biomedical (Lee et al., 2019; Gu et al., 2021; Luo et al., 2022; Singhal et al., 2023a, b) and legal (Chalkidis et al., 2019; Zheng et al., 2021; Henderson et al., 2022; T.y.s.s et al., 2024; Niklaus et al., 2024) domains. See Appendix C for task examples.
@@
We use the MultiMedQA benchmark (Singhal et al., 2023a), specifically the PubMedQA (Jin et al., 2019), MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022) tasks, and relevant subsets from MMLU (Hendrycks et al., 2021) (anatomy, clinical knowledge, college medicine, medical genetics, professional medicine, college biology). To assess language modeling performance, we use the BioLaySumm PLOS dataset (Goldsack et al., 2022, 2023) comprising biomedical articles.
@@ 5.1 Pruning
Figure 3 presents the benchmark accuracy when pruning to 50% sparsity, relative to the original (dense) model. [^2] The D-Pruner (Zhang et al., 2024) implementation only supports models with the Llama architecture. See Limitations. We report the mean value and standard deviation across five calibration sets. For brevity, we present the general performance across domain-specific models. We additionally present complete results across all models in Appendix E.
@@
To maximize the performance of the D-Pruner baseline, we perform an extensive hyperparameter search across $\lambda\in\{0.1,0.01,0.001\}$ and group size $\in\{\text{None},128\}$ for each model and domain. We then present results for only the best performing combinations. We present complete results across all hyperparameters in Appendix E. In contrast, we do not optimize $\alpha$ for cross-calibration (ours) and simply opt to use $\alpha=0.8$ across all models and domains. We ablate the impact of this hyperparameter in Appendix A.3.
Cross-calibration benefits domain performance.
We observe that across both biomedical and legal domains, our approach consistently outperforms all other compression methods. For example, we observe that cross-calibration achieves an average relative accuracy of 91.3% on the legal benchmark for Llama 3.1 8B. In comparison, SparseGPT and D-Pruner achieve 72.9% and 77.2%, respectively. For the biomedical benchmark, a similar trend can be observed. Cross-calibration achieves 88.4%, while SparseGPT and D-Pruner achieve 81.0% and 75.3%, respectively. This highlights that cross-calibration is effective at isolating weights that are influential for strong in-domain performance.
Figure 4: Average accuracy when applying 4-bit quantization and 2:4 sparsity, relative to the dense model.
@@
In addition to substantial improvements in domain-specific performance, general performance remains similar to the general-purpose pruning baselines. With Llama 3.1 8B, cross-calibration achieves 91.8% relative accuracy for general tasks on average. In comparison, SparseGPT achieves a slightly higher value of 92.2%, while D-Pruner reaches only 85.7%. Intriguingly, for the Gemma family of models, we note that cross-calibration achieves higher performance than even the general-purpose methods. For example, cross-calibration achieves 94.3% on average with Gemma 2 9B, while SparseGPT reaches 93.7%. This suggests that cross-calibration is able to effectively incorporate domain-specific features without sacrificing general performance.
@@
Finally, we observe that the performance benefits of cross-calibration are similar irrespective of the model size and family. For example, we consider two similarly sized models, Llama 3.1 8B and Gemma 2 9B. In the biomedical benchmark, we observe a 13.1 and 8.7 point increase in relative accuracy over SparseGPT, respectively. For the legal benchmark, we observe an 18.4 and 16.3 point increase, respectively. Our findings indicate that cross-calibration and the resulting performance gains are agnostic to the model size and family.
@@
Table 7 (Appendix E) presents perplexity results. Similar to the downstream task experiments, we note that cross-calibration achieves the best performance on in-domain language modeling. Using Llama 3.1 8B, cross-calibration achieves a perplexity of 5.9 compared to 6.7 from D-Pruner in the legal domain. For the biomedical domain, cross-calibration has a perplexity of 11.0, compared to 15.1 from D-Pruner. The datasets used to evaluate perplexity are not used for calibration, suggesting that cross-calibration succeeds at identifying domain-specific features.
Cross-calibration appears to generalize beyond specific tasks.
Across all models and domains, we observe that the performance benefits of cross-calibration continue to tasks not included in the calibration data. In Table 10 (Appendix E), we present complete per-task results. We consider the MMLU tasks, which notably do not have training data, and are therefore not represented in the calibration data. For Llama 3.1 8B, cross-calibration achieves an absolute increase in accuracy of 8.6 points over SparseGPT in the biomedical domain. In the legal domain, cross-calibration achieves an additional 3.3 points in accuracy. This suggests that it identifies features that are domain-relevant, rather than relevant to only a specific task.
@@ 5.2 Joint Pruning and Quantization
We further examine the performance of our approach when jointly applying pruning and quantization by reusing the same inverse Hessian (Frantar and Alistarh, 2023). This has the advantage of allowing pruning and quantization decisions to influence each other and enables quantization at almost no extra cost. Figure 4 presents benchmark accuracy when jointly applying 2:4 sparsity with 4-bit quantization, relative to the original model.
Cross-calibration improves in-domain performance while sustaining general performance.
Cross-calibration achieves substantially greater domain performance than SparseGPT for 2:4 sparsity. For example, it sees a relative accuracy of 79.9% on the biomedical benchmark for Gemma 2 9B, versus 58.2% with SparseGPT. For general performance, we observe a similar trend to the pruning results, with cross-calibration performing comparably to SparseGPT. This illustrates that cross-calibration can be reliably used with quantization.
@@ 5.3 Compression Efficiency
Cross-calibration does not sacrifice efficiency.
Figure 5 presents the time and memory requirements of each method, which are often limiting factors in practice. First, we observe that cross-calibration does not increase the duration of compression over SparseGPT, both at 0.8 hours. This is considerably faster than D-Pruner, which on average takes 9.3 hours, 11 times longer than our approach. We also observe that cross-calibration does not increase the memory required for compression over SparseGPT, with both using up to 20 GB of memory. In contrast, D-Pruner uses up to 65.6 GB of memory, over three times more than our approach. Consequently, this suggests that cross-calibration is more practical than D-Pruner, with lower computational requirements.
@@ 6 Analysis
6.1 Ablating Cross-calibration
To better understand the role that in-domain data plays in pruning, we conduct an ablation study on cross-calibration. Table 1 presents the performance impact versus a SparseGPT baseline. Namely, we examine the in-domain and general performance when (1) employing domain-specific calibration data, and (2) incorporating a mixture of Hessians (i.e. cross-calibration).
Domain-specific data benefits in-domain performance, however harms general performance.
We observe that using domain-specific calibration data can substantially improve in-domain performance. For example, Llama 3.1 8B sees an average increase in accuracy of 8.1 and 10.2 points for the biomedical and legal domains, respectively. However, these improvements come at the cost of general performance. For example, general performance drops by 1.0 and 2.2 points when employing biomedical and legal calibration data, respectively.
Hessian mixing maximizes overall performance.
We observe that by using a mixture of Hessians, overall performance is higher than using domain-specific calibration data alone. This comes at almost no cost to in-domain performance, with a negligible reduction in both the biomedical and legal domains. This suggests that the addition of Hessian mixing can enable an optimal balance between domain-specific and general performance.
6.2 Performance in Other Languages
To explore whether cross-calibration generalizes beyond English, we further experiment with a Chinese-language model and benchmarks. We select the Chinese language as it is morphologically distinct from English and well-resourced in terms of models and domain-specific evaluation tasks.
@@ 6.2 Performance in Other Languages
SparseGPT + CC (Ours) 53.70.5 63.00.3 58.40.3
-- Hessian Mixing 53.90.5 61.60.3 57.70.3
-- In-domain Data 45.81.4 62.60.2 54.20.7
@@
SparseGPT + CC (Ours) 50.20.4 61.60.2 55.90.2
-- Hessian Mixing 50.30.3 60.40.4 55.30.1
-- In-domain Data 40.10.6 62.60.2 51.30.4
Table 1: Average benchmark accuracy when pruning Llama 3.1 8B. Ablations are presented cumulatively.
@@
Cross-calibration appears language-agnostic.
Figure 6 presents the relative benchmark accuracy when pruning to 50% sparsity, similar to earlier experiments (§5.1). The results indicate that cross-calibration outperforms other pruning approaches for domain performance, while maintaining comparable general performance to SparseGPT. For example, cross-calibration achieves a relative accuracy of 86.8% in the biomedical domain, compared to 70.0% and 79.6% from SparseGPT and D-Pruner, respectively. In the legal domain, the performance retention from cross-calibration (92.7%) is substantially greater than both SparseGPT (47.1%) and D-Pruner (76.2%). These findings are in-line with the English language tasks, indicating that in-domain performance benefits from cross-calibration transfer to the Chinese language. This indicates that our proposed approach is language agnostic.
Figure 6: The mean accuracy for Yi 1.5 6B on Chinese benchmarks, relative to the original model.
@@ 7 Conclusion
In this paper, we proposed cross-calibration as a solution for creating compressed LMs for specialized domains. We empirically validated cross-calibration using a plethora of pre-trained models and evaluation tasks. Our approach represents a substantial advancement over earlier work such as D-Pruner, offering consistent performance improvements with a smaller computational footprint. We hope that our study will inspire further work towards the efficient deployment of LMs in specialized domains. As future work, we are interested in exploring the role that continual pre-training could play in further enhancing the performance of compressed domain-specific LMs (Cheng et al., 2024; Xie et al., 2024).
@@
The D-Pruner (Zhang et al., 2024) implementation supports only the Llama model architecture. [^3] https://github.com/psunlpgroup/D-Pruner Consequently, we are limited to offering comparisons for only the models using the Llama architecture (i.e. Llama 3.2 3B, Yi 1.5 6B and, Llama 3.1 8B). We emphasize that our approach substantially outperforms D-Pruner across all tested models and domains. Therefore, we expect that this trend would continue for the models not supported by D-Pruner.
@@
Our work enables the efficient and effective compression of LMs for specialized domains. We note that this poses dual-use concerns, as it may enable misuse at a lower cost (Weidinger et al., 2022). However, we emphasize that our approach is unlikely to enhance or introduce new harmful abilities, as the performance of compressed models is constrained by the the capabilities of the original.
@@ Appendix A Additional Ablation Studies
We further examine the performance and transferability of our approach when using quantization. Figure 4 presents the benchmark accuracy when applying 4-bit quantization with GPTQ-M and GPTQ-M + CC (Ours), relative to the original model.
Cross-calibration generalizes to quantization.
Figure 7 (Appendix A.1) shows the performance of cross-calibration in quantization. This uncovers the extent to which cross-calibration benefits each method independently in the joint sparsification and quantization experiments (§5.2). We find that cross-calibration performs better than vanilla GPTQ-M for in-domain performance across the board (4 out of 5 in the legal domain and 5 out of 5 models in the biomedical domain). General performance remains also unaffected, with cross-calibration performing equally or better than GPTQ-M across all models. This suggests that the benefits to in-domain performance are transferable across Hessian-based compression methods.
Cross-calibration improves domain performance, yet sustains general performance.
We observe that cross-calibration can achieve greater domain performance than GPTQ-M alone. For example, cross-calibration sees a relative accuracy of 97.4% on the biomedical benchmark for Llama-3.1 8B, versus 94.7% with GPTQ-M. We note that the performance improvements for quantization are smaller than in the pruning experiments. This is expected, as quantization is less sensitive to calibration data (Williams and Aletras, 2024). Considering general performance, we observe a similar trend to the pruning results, with cross-calibration performing comparably to GTPQ-M. This illustrates that cross-calibration can be reliably used in quantization.
@@ A.2 Performance Across Sparsity Levels
In Figure 8, we examine how cross-calibration (Ours) performs comparatively to the other pruning methods across different sparsity levels. We report the average benchmark accuracy (legal and biomedical) with standard deviation for each sparsity levels shown by the shaded regions.
We first observe that in-domain performance with cross-calibration remains consistently higher than other approaches even beyond 50% sparsity. For example, the legal benchmark accuracy at 60% sparsity is at 46% with cross-calibration compared to less than 30% with SparseGPT. Similarly, performance in generic reasoning tasks remains comparable to SparseGPT up to 90% sparsity levels. This suggests that cross-calibration can effectively isolate domain-specific features without sacrificing generic performance.
A.3 Cross-calibration Hyperparameter
In Figure 9, we examine how different regularization of the Hessian (by modifying $\alpha$) impacts in-domain and general performance at 50% unstructured sparsity. The relative improvement over SparseGPT is presented, with dashed lines representing the general performance. The continuous lines represent biomedical (Figure 9(a)) and legal (Figure 9(b)) performance.
Figure 7: Average accuracy when applying 4-bit quantization, relative to the dense model.
First, we observe that in-domain performance (i.e. biomedical or legal) increases by increasing the contribution of the in-domain computed Hessian. Considering Gemma 2 2B in the biomedical domain, we observe that $\alpha=0.1$ results in a 1.25$\times$ relative improvement compared to SparseGPT, whilst $\alpha=0.9$ leads to approximately a 1.30$\times$ relative improvement in average task accuracy. Surprisingly, we find that even a small contribution from the in-domain Hessian ($\alpha=0.1$) is enough to offer improvements for in-domain performance. For example, in the Legal domain with Llama 3.2 3B and $\alpha=0.1$ we observe a relative improvement of approximately 1.14$\times$ .
Reasoning performance remains relatively consistent, until $\mathbf{\boldsymbol{\alpha}=1}$.
We first observe that reasoning performance for all models (dashed lines), largely remains competitive with $\alpha=0$ across almost all balance levels. Our findings translate across all models tested, with some models benefiting more than others. On average, we found $\alpha=0.7$ to strike a good balance between in-domain and general performance. This suggests that $\alpha$ could be treated as an optimizable hyper-parameter, however this is not essential.

### v2

@@ 4.2 Domains and Tasks
To assess the efficacy of our approach on downstream tasks, we experiment with two of the most widely explored domains in NLP, the biomedical (Lee et al., 2019; Gu et al., 2021; Luo et al., 2022; Singhal et al., 2023, 2025) and legal (Chalkidis et al., 2019; Zheng et al., 2021; Henderson et al., 2022; T.y.s.s et al., 2024; Niklaus et al., 2024) domains. See Appendix C for concrete task examples.
@@
We use the MultiMedQA benchmark (Singhal et al., 2023), specifically the PubMedQA (Jin et al., 2019), MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022) tasks, and relevant subsets from MMLU (Hendrycks et al., 2021) (anatomy, clinical knowledge, college medicine, medical genetics, professional medicine, college biology). To assess language modeling performance, we use the BioLaySumm PLOS dataset (Goldsack et al., 2022, 2023) comprising biomedical articles.
@@ 5.1 Pruning
Figure 3 presents the benchmark accuracy when pruning to 50% sparsity, relative to the original model. [^2] The D-Pruner (Zhang et al., 2024) implementation only supports models with the Llama architecture. See Limitations. We report the mean value and standard deviation across five calibration sets. For brevity, we present the average performance across domain-specific models. We additionally present complete results across all models in Appendix E.
@@
To maximize the performance of the D-Pruner baseline, we perform an extensive hyperparameter search across $\lambda_{g}\in\{0.1,0.01,0.001\}$ and group size $\in\{\text{None},128\}$ for each model and domain. We then present results for only the best performing combinations. We present complete results across all hyperparameters in Appendix E. In contrast, we do not optimize the hyperparameter for our approach ($\alpha$) and simply use 0.5 across all models and domains. We ablate the impact of this hyperparameter in Appendix A.3.
MixCal benefits domain performance.
We observe that across both biomedical and legal domains, our approach consistently outperforms all other compression methods. For example, we observe that MixCal achieves an average relative accuracy of 88.6% on the legal benchmark for Llama 3.1 8B. In comparison, SparseGPT and D-Pruner see 72.9% and 77.2%, respectively. For the biomedical benchmark, a similar trend can be observed. MixCal achieves 86.5%, while SparseGPT and D-Pruner reach 75.3% and 81.0%, respectively. This indicates that MixCal can be effective at identifying weights influential to in-domain performance.
Figure 4: Average accuracy when applying 4-bit quantization and 2:4 sparsity, relative to the original model.
@@
In addition to substantial improvements on domain-specific benchmarks, general performance remains similar to the general-purpose pruning baselines. With Llama 3.1 8B, MixCal achieves 92.6% relative accuracy for general tasks on average. In comparison, SparseGPT achieves a slightly lower value of 92.2%, while D-Pruner reaches only 85.7%. Intriguingly, we note that MixCal achieves higher performance than even the general-purpose methods for most model families. For example, MixCal sees an average of 94.5% with Gemma 2 9B, while SparseGPT reaches 93.7%. This suggests that MixCal can effectively establish domain-specific features without sacrificing general performance.
@@
Finally, we observe that the performance benefits of MixCal are similar irrespective of the model size and family. For example, we consider two similarly sized models, Llama 3.1 8B and Gemma 2 9B. In the biomedical benchmark, we observe an 11.2 and 7.5 point increase in relative accuracy over SparseGPT, respectively. For the legal benchmark, we see a 15.7 and 16.1 point increase. This suggests that the performance gains from MixCal are independent of the model family.
@@
Table 7 (Appendix E) presents perplexity results. Similar to the downstream task experiments, we note that MixCal achieves the best performance on in-domain language modeling. Considering Llama 3.1 8B, MixCal achieves a perplexity of 6.0 compared to 6.7 with D-Pruner for the legal domain. For the biomedical domain, MixCal has a perplexity of 10.7, compared to 14.7 from D-Pruner. The datasets used to evaluate perplexity are not used for calibration, suggesting that MixCal assists with identifying domain-specific features.
MixCal appears to generalize beyond specific tasks.
Across all models and domains, we observe that the performance benefits of MixCal continue to tasks not included in the calibration data. In Table 10 (Appendix E), we present complete per-task results. We consider the MMLU tasks, which do not have training data, and are therefore not represented in the calibration data. For Llama 3.1 8B, MixCal achieves an absolute increase in accuracy of 7.7 points over SparseGPT in the biomedical domain. In the legal domain, MixCal achieves an increase of 2.7 points in accuracy. This suggests that MixCal can identify features that are relevant to the domain, rather than only a specific task.
@@ 5.2 Joint Pruning and Quantization
We further examine the performance of our approach when jointly applying pruning and quantization by reusing the same inverse Hessian (Frantar and Alistarh, 2023). This has the advantage of allowing pruning and quantization decisions to influence each other, and enables quantization at almost no extra cost. Figure 4 presents benchmark accuracy when jointly applying 2:4 sparsity with 4-bit quantization, relative to the original model.
MixCal improves in-domain performance while sustaining general performance.
MixCal achieves substantially greater domain performance than SparseGPT for 2:4 sparsity. For example, it sees a relative accuracy of 77.8% on the biomedical benchmark for Gemma 2 9B, versus 58.2% with SparseGPT. For general performance, we observe a similar trend to the pruning results, with MixCal also performing comparably. This illustrates that MixCal can be reliably used with quantization.
@@ 5.3 Compression Efficiency
MixCal does not sacrifice efficiency.
Figure 5 presents the time and memory requirements of each method, which are often limiting factors in practice. First, we observe that MixCal (0.5 hours) reduces the duration of compression compared to SparseGPT (0.8 hours). This is considerably faster than D-Pruner, which on average takes 9.3 hours, 18 times longer than our approach. We also observe that MixCal does not increase the memory required for compression over SparseGPT, with both using up to 20 GB of memory. In contrast, D-Pruner uses up to 65.6 GB of memory, over three times more than our approach. This suggests that MixCal is more practical than D-Pruner, offering lower computational resource requirements.
@@ 6 Analysis
6.1 Ablating MixCal
To better understand the role that in-domain data plays in pruning, we conduct an ablation study of MixCal. Table 1 presents the in-domain and general language modeling performance (perplexity) versus a SparseGPT baseline with generic calibration data. We specifically show the effect of (1) adding in-domain data to the calibration data mixture, and (2) additionally using MixCal. In both cases, we use the same equal mixture of in-domain and generic data. Results are averaged across models, with granular results in Table 11, Appendix E.
@@ 6.1 Ablating MixCal
SparseGPT 16.4 14.0 15.2
+ In-domain data 15.1 13.8 14.4
+ MixCal (Ours) 14.4 13.4 13.9
@@
SparseGPT 7.2 14.0 10.6
+ In-domain data 6.7 13.7 10.2
+ MixCal (Ours) 6.7 13.5 10.1
Table 1: Perplexity when pruning to 50% unstructured sparsity with (1) in-domain data, and (2) MixCal, averaged across all models.
Domain-specific data benefits language modeling performance.
We first observe that using a mix of domain-specific and generic calibration data can substantially improve in-domain performance. For example, when using a mix of in-domain data, we find lower perplexity scores for both domains. Specifically, we observe an average perplexity of 15.1 versus 16.4 for the biomedical domain, and 6.7 versus 7.2 for the legal domain. This highlights the influential role played by the calibration data, corroborating Williams and Aletras (2024).
MixCal maximizes overall performance.
Finally, we observe that by applying MixCal, language modeling performance is maximized compared to introducing domain-specific calibration data alone. For example, MixCal achieves a reduction of 0.7 (average) perplexity for the biomedical domain (14.4 versus 15.1). For the legal domain, we note that the average perplexity is equivalent. Considering overall performance, i.e. the average of domain-specific and general performance, MixCal yields the best results. This suggests that the addition of MixCal can balance domain-specific and general performance effectively.
6.2 Performance in Other Languages
To explore whether MixCal generalizes beyond English, we further experiment with a Chinese-language model and benchmarks. We select the Chinese language as it is morphologically distinct from English and well-resourced in terms of models and domain-specific evaluation tasks.
@@ 6.2 Performance in Other Languages
Figure 6: The mean accuracy for Yi 1.5 6B on Chinese-language benchmarks, relative to the original model.
MixCal appears language-agnostic.
Figure 6 presents the relative benchmark accuracy when pruning to 50% sparsity, similar to earlier experiments (§5.1). The results indicate that MixCal outperforms other pruning approaches for domain performance, while maintaining comparable general performance to SparseGPT. For example, MixCal achieves a relative accuracy of 85.6% in the biomedical domain, compared to 70.0% and 79.6% from SparseGPT and D-Pruner, respectively. In the legal domain, the performance retention from MixCal (90.1%) is substantially greater than both SparseGPT (47.1%) and D-Pruner (76.2%). These findings are in line with the English language tasks, suggesting that MixCal is language-agnostic.
@@ 7 Conclusion
In this paper, we proposed MixCal as a solution for creating compressed LMs for specialized domains. We empirically validated MixCal using a plethora of pre-trained models and evaluation tasks. Our approach represents a substantial advancement over earlier work such as D-Pruner, offering consistent performance improvements with a smaller computational footprint. We hope that our study will inspire further work towards the efficient deployment of LMs in specialized domains. As future work, we are interested in exploring how synthetic calibration data could be incorporated to further enhance the performance of domain-specific LM compression (Williams et al., 2025).
@@
The D-Pruner (Zhang et al., 2024) implementation supports only the Llama model architecture. [^3] https://github.com/psunlpgroup/D-Pruner Consequently, we are limited to offering comparisons for only the models using the Llama architecture (i.e. Llama 3.2 3B, Yi 1.5 6B, and Llama 3.1 8B). We emphasize that our approach substantially outperforms D-Pruner across all tested models and domains. Therefore, we expect that this trend would continue for the models not supported by D-Pruner.
@@
Our work enables the efficient and effective compression of LMs for specialized domains. We note that this poses dual-use concerns, as it may enable misuse at a lower cost (Weidinger et al., 2022). However, we emphasize that our approach is unlikely to enhance or introduce new harmful abilities, as the performance of compressed models is constrained by the capabilities of the original.
Acknowledgments
We would like to thank Huiyin Xue and the anonymous reviewers for their invaluable feedback. Additionally, we are sincerely grateful to Huiyin Xue for assistance with the Chinese-language tasks. MW is supported by the Centre for Doctoral Training in Speech and Language Technologies (SLT) and their Applications funded by UK Research and Innovation grant EP/S023062/1. NA is supported by EPSRC grant EP/Y009800/1, part of the RAI UK Keystone projects.
@@ Appendix A Additional Ablation Studies
MixCal generalizes to quantization.
We further examine the performance and transferability of our approach under the setting of quantization. Figure 7 presents the benchmark accuracy when applying 4-bit quantization with GPTQ-M and MixCal, relative to the original model. This uncovers the extent to which MixCal benefits each method independently in the joint sparsification and quantization experiments (§5.2). We observe that general performance is consistently preserved, with GPTQ-M + MixCal performing comparably to GPTQ-M across all models. This suggests that MixCal is transferable across settings in Hessian-based compression.
MixCal improves in-domain performance.
We observe that MixCal can achieve greater domain performance than GPTQ-M alone. For example, MixCal sees a relative accuracy of 96.6% on the biomedical benchmark for Llama 3.1 8B, versus 94.7% with GPTQ-M (Figure 7). This indicates that MixCal can be reliably used for LM quantization. We note that the performance gains for quantization are smaller than in the pruning experiments. This is expected, as quantization is less sensitive to calibration data at the tested sparsity level and bit width (Williams and Aletras, 2024).
@@ A.2 Performance Across Sparsity Levels
In Figure 8, we examine how MixCal performs compared to the other pruning methods across different sparsity levels. We report the average in-domain and general benchmark performance, with standard deviation denoted by the shaded regions. We first observe that in-domain performance with MixCal consistently remains higher than other approaches, even beyond 50% sparsity. For example, accuracy on the legal benchmark at 60% sparsity is 43.3% with MixCal, compared to 27.3% with SparseGPT. Similarly, general benchmark performance remains comparable to SparseGPT across all sparsity levels. This suggests that MixCal can be used to effectively isolate domain-specific features without sacrificing generic performance.
A.3 MixCal Hyperparameter Analysis
Figure 7: Average accuracy when applying 4-bit quantization, relative to the original model.